1. What will be the elasticity of resource demand in the following cases? Describe each case. (6 marks)
  2. unit wages rise by 10 percent and the number of employed workers falls by 5 percent.
  3. unit wages rise by 4 percent andthe number of employed workers falls by 6 percent.
  4. unit wages rise by 3 percent and thenumber of employed workers falls by 3 percent.
1
Expert's answer
2021-06-16T09:26:45-0400

2."\frac {5}{10}=0.5"


3."\frac {6}{4}=1.5"


4."\frac {3}{3}=1"